4. You are the manager of a monopoly, and your demand and cost functions are given by P-500- Q and C(Q)- 6,000+402, respectively. A. B. What price-quantity combination maximizes your firms profits? Calculate the maximum profits.
Answer #1

a) Given that demand function is P = 500 - Q and C = 6000 + 4Q^2

Profit = revenue - cost

= PQ - C

= (500 - Q)Q - 6000 - 4Q^2

= 500Q - 6000 - 5Q^2

Profits are maximized when marginal profits are zero

500 - 10Q = 0

Q = 50 units

P = 500 - 50 = $450

Hence profit maximizing price is $450 per unit and quantity is Q = 50 units

b) Profits are maximized at = 500x50 - 6000 - 5x(50^2) = $6500.